OpenStudy (anonymous):

Evaluate 3! + 4! can someone explain me the steps!

OpenStudy (anonymous):

By order of operations, you'll want to evaluate the factorials first. So calculate 3! and 4! and then add the two numbers together

OpenStudy (anonymous):

but the answers choices are 30, 144, 5040!:(((

OpenStudy (anonymous):

Haha well, the right answer is there... Let's try this: 3! = 3 x 2 x 1 = ? 4! = 4 x 3 x 2 x 1 OR 4! = 4 x 3! = ?

OpenStudy (anonymous):

really :/ having a hard time! not understanding the steps...

OpenStudy (anonymous):

Do you understand what the factorial operation (!) does?

OpenStudy (anonymous):

no....:/

OpenStudy (anonymous):

No problem :) Let me explain So the exclamation point is called a factorial. Basically what it does is, say you see 5!, then you multiply all the numbers starting from 5 down to 1. Like, 5! = 5 * 4 * 3 * 2 * 1 Similarly, for 4!, when you count from 4 to 1 you hit 4, 3, 2, 1. So 4! = 4 * 3 * 2 * 1 Does that make sense?

OpenStudy (anonymous):

would the answer be 25???

OpenStudy (anonymous):

For which question?

OpenStudy (anonymous):

4!=4*3*2*1

OpenStudy (anonymous):

You're close. 4*3 is 12, so 4*3*2 = 12*2 = 24!

OpenStudy (anonymous):

you wouldn't add the one???

OpenStudy (anonymous):

This is multiplication. You do multiply the 1, but remember that 1 times any number is just that number. so 24*1 = 24

OpenStudy (anonymous):

3!+4! would it be 12 *7???

OpenStudy (anonymous):

3! = 3*2*1 4! = 4*3*2*1 So 3! + 4! = 3*2*1 + 4*3*2*1 = 6 + 24 Does that make sense?

OpenStudy (anonymous):

so the answer is 30??

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thankyou so much!!!:))))